club A offers membership for a fee of 30$ plus a monthly foo of 25$. Club B offer membership for a fee of 20$ plus a monthly fee of 15$. After how many months will the total cost at each health club be the same? what will be the total cost for each club?
Answer by Boreal(15235)

A: $30 +$25 x, where x=number of months
B: $20 + $15 x
Set them equal
30 +25x = 20 +15 x
They will never be equal as the problem is stated. The second one costs less to start and has fewer monthly charges. Let me switch the monthly charges.
30 + 15x= 20 +25x
This will be equal at one month at $45.